OpenStudy (anonymous):

how do you integrate X/(sqrt(1-x^2)

OpenStudy (anonymous):

use u-substitute. let u = sqrt (1-x^2) u^2 = 1 - x^2 2udu = 1 - 2xdx since your numerator is just xdx...you isolate xdx 2xdx = 1 - 2udu xdx = (1-2udu)/2 i assume you know the rest?

OpenStudy (anonymous):

i'm actually helping out a friend who's in calc and i haven't done this for so long! could you please help me finish out the problem?

OpenStudy (anonymous):

since xdx = (1-2udu)/2 and sqrt(1-x^2) is u then your integral becomes \[\huge \int\frac{\frac{1-2udu}2}u\] right?

OpenStudy (anonymous):

okay got that

OpenStudy (anonymous):

now take out that over two since it's just a coefficient \[\huge \frac 12\int \frac{1-2udu}u\]

OpenStudy (anonymous):

then since the numerator is subtraction, you can separate the integrals \[\huge \frac 12 [\int \frac 1u du - \int \frac{2udu}u]\] do you now know the rest?

OpenStudy (anonymous):

yeah i got it from there! thanks so much for the help!

OpenStudy (anonymous):

You also could have made the substitution \[u=1-x^2~\Rightarrow~-\frac{1}{2}du=x~dx\] (same thing as @roxygurl453's suggestion, really) or a trigonometric substitution.

OpenStudy (anonymous):

Sorry, meant @abb50!

OpenStudy (anonymous):

ohhh okay haha thanks ! i appreciate the help a lot !!!!